slave618:

A football quarterback enjoys practicing his long passes over 40 yards. He misses the first pass 40% of the time. When he misses on the first pass, he misses the second pass 20% of the time. What is the probability of missing two passes in a row?

Florisalreadytaken:

ok so the probability for missing the first shot is 40% -- then, the probability for missing the 2nd shot is 20% -- one after the other -- this would give us a conditional probability. lets say that the prob. for missing the first shot will be \( P(1) \), and the probability for missing the 2nd shot will be \( P(2) \) \[ P(1 \cap 2) = P(1) \times P(2) \]

Florisalreadytaken:

lets turn the percentages in decimal form \( 40\%= \frac{40}{100} \) \( 12\%= \frac{20}{100} \)

Florisalreadytaken:

\( 40\%= 0.4 \) \( 20\%=0.2 \) thus, \[ P(1 \cap 2) = 0.4 \times 0.2 \] do the math, and that would be our answer!

Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment
Sovereign:

1 attachment